OpenStudy (anonymous):

4(9+3t)-12=-6 What is t?

OpenStudy (anonymous):

36+12t-12=-6 24+12t=-6 12t=-30 t = -30/12 t = -5/2

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

lol:)